# Grounding Text-To-Image Diffusion Models For Controlled High-Quality Image Generation

### v1

@@ 4.4 Limitations and Failures
Table 4: This table lists the quality evaluation results of generated images from COCO2017 [30] validation set. We use FID [17] score implementation from [55] and resize the real and generated images to 299×299 before calculating the FID score. Other FID scores are taken from [28].
@@ 5 Discussion and Conclusion
In this work, we have taken inspiration from the two prominent works, ControlNet and GILGEN to propose a new model ObjectDiffusion. Our model learns to ground T2I models on entity name and bounding box annotations. ObjectDiffusion exceeds the current state-of-the-art GLIGEN (fine-tuned) trained on open-source datasets in AP50 and AR metrics. Our model also scores higher than the top-scoring layout-to-image baselines. In the FID quality metric, our model surpasses the top scores of both specialized layout-to-image and diffusion-based models conditioned on COCO2017 object detection annotations. Qualitative evaluation demonstrates the ability of ObjectDiffusion to synthesize diverse high-quality, high-fidelity images that seamlessly adhere to the conditional inputs.
While ObjectDiffusion has demonstrated a remarkable performance, several improvements may further enhance the model’s capabilities. First, training on a large-scale dataset that contains precise and detailed entity names could bring a significant increase in AP score. Second, further training our model may lead to a pronounced gain in the overall performance. Additionally, training on powerful GPUs would enable fitting a larger batch size of 32 or 64 and allows the use of full float32 precision. While we achieve a remarkable improvement, scaling up the training has the potential to accelerate the model convergence and further improve the generation precision and quality.
@@ Acknowledgment
Figure 6: This figure represents qualitative examples of closed-set configuration. Images are generated from COCO2017 [30] validation set annotations. Our model demonstrates precise generation capabilities of complex indoor visuals, for example, living rooms, kitchens, and bathrooms.
@@
Figure 7: This figure represents qualitative examples of closed-set configuration. Images are generated using grounding entities and image captions from COCO2017 [30] validation set annotations. Our framework can accurately render different classes of vehicles like cars, trains, motorcycles, and boats.
@@
Figure 8: This figure represents qualitative examples of closed-set configuration. Images are generated using grounding entities and image captions from COCO2017 [30] validation set annotations. The portraits generated by ObjectDiffusion feature realistic facial expressions.
@@
Figure 9: This figure represents qualitative examples of closed-set configuration. Images are generated using grounding entities and image captions from COCO2017 [30] validation set annotations. Our framework achieves proficiency in generating high-resolution visuals of animals in their native habitat with a very high AP precision score.
@@
Figure 10: This figure represents qualitative examples of closed-set configuration. Images are generated using grounding entities and image captions from COCO2017 [30] validation set annotations. This figure shows static and dynamic scenes of different object categories. ObjectDiffusion can synthesize images of dishes and visuals of sport activities, for example, surfing and skiing
(a) Image Caption: “bright modern Scandinavian style house with a comfy couch, stylish coffee table, and sunlit windows homevibes relaxation” Conditional Entities: a comfy couch, a stylish coffee table, a sunlit windows
(b) Image Caption: “a sleek, ultra-thin, high resolution bezel-less monitor, realistic, modern, detailed, vibrant colors, glossy finish, floating design, lowlight, art by peter mohrbacher and donato giancola, digital illustration, trending on Artstation, high-tech, smooth, minimalist workstation background, crisp reflection on screen, soft lighting” Conditional Entities: a sleek, ultra-thin, high resolution bezel-less monitor, a keyboard, a mouse, a studying desk, a vase, a window
(c) Image Caption: “a waterfall and a modern high speed train running through the tunnel in a beautiful forest with fall foliage” Conditional Entities: a waterfall, a modern high speed train
(d) Image Caption: “a beautiful digital painting depecting fluffy Pomeranian in a green forest, impressionism painting style, impressive, masterpeace” Conditional Entities: a fluffy Pomeranian
Figure 11: This figure represents realistic and artistic qualitative examples of open-set configuration with manual annotations.

### v2

@@ 4.3 Qualitative Results
Table 4: This table lists the quality evaluation results of generated images from the COCO2017 [30] validation set. ObjectDiffusion achieves the highest quality, setting a new benchmark in conditional image generation. We use the FID [17] score implementation from [56] and resize both the real and generated images to 299 $\times$ 299 before calculating the FID score. Other FID scores are taken from [28].
The first row (a) demonstrates image examples of a modern house in Scandinavian style. We can use the category names from the 80 COCO categories but with modifier attributes preceding the name, such as comfy couch, stylish coffee table, or we can use completely new concepts, such as sunlit windows. We observe a noticeable improvement in the quality of images in the open-set vocabulary setting compared to their counterparts in the closed-set vocabulary setting. This enhancement may be explained by the flexibility that the open-ended text provides.
Second row (b) depicts visuals of fancy desk setups. ObjectDiffusion can understand long and complex captions and entity texts, such as a sleek, ultra-thin, high-resolution bezel-less monitor. Also, our model was creative in rendering the three vases, each with a distinct shape.
Third row (c) contains stunning images of a modern high-speed train passing beside a waterfall. Even though the waterfall is an uncommon and has never been seen as an instance name in the COCO2017 dataset, our model is able to interpret and generate it. However, we observe that all of our output images miss the tunnel. This may be because it was not in the grounding entity input. The open-set qualitative visuals indicate that ObjectDiffusion is able to retain the knowledge of the pretrained SD [50] and GLIGEN [28] models while successfully demonstrating conditioning capabilities on the fine-tuning grounding task.
The last row (d) in Figure 11 emphasizes the powerful regional grounding skills of our model. The layout is constructed by stacking three bounding boxes of varying sizes, all belonging to the same object name, a fluffy Pomeranian. As seen in the synthesized images, ObjectDiffusion succeeds in generating the correct entity with the correct size in the correct location. Furthermore, our model exhibits a high level of diversity, rendering the Pomeranians in three different colors: brown, white, and black, and with three different background illumination contrasts.
Generated images from the open-set vocabulary settings demonstrate the capabilities ObjectDiffusion of synthesizing realistic and artistic visuals, both of which are spectacular, diverse, and high-quality.
4.4 Limitations and Failures
ObjectDiffusion is able to synthesize high-quality scenes; however, it sometimes struggles with capturing the intricate fine-grained details of faces, hands, and feet [52]. These are highly detailed, complex structures that exhibit a high degree of variability. Image (a) in Figure 5 is an example of a young man with a distorted face and hands.
Similar to the limitation in face synthesis, ObjectDiffusion exhibits an imperfection in rendering images that contain text. We observe that our model repeatedly fails to render images with embedded text. As seen in image (b) in Figure 5, our model fails to overlay a short, single-word text on the stop sign, even though our fine-tuning dataset contains an abundance of stop sign images.
ObjectDiffusion may fail to account for the spatial relationships and interactions between objects in the scene. This leads to unrealistic object placements. For instance, in image (c) of Figure 5, while the dog and the vehicle are correctly positioned within their designated bounding boxes, the legs of the dog do not make contact with the truck, creating an unnatural scene where the dog seems to float, disconnected from its surroundings.
@@ 5 Discussion and Conclusion
[…]
While ObjectDiffusion has demonstrated remarkable performance, several potential improvements could further enhance the capabilities of our model. First, training on a large-scale dataset containing precise and detailed entity names could notably increase the AP score. Second, training our model for more iterations may lead to a noticeable improvement in overall performance. Additionally, training on powerful GPUs would allow for a larger batch size and enable the use of full precision. Scaling up the training has the potential to accelerate model convergence and further improve generation precision and quality.
@@ Acknowledgment
Figure 6: This figure represents qualitative examples of closed-set configuration. Our model demonstrates precise generation capabilities for complex indoor visuals, such as living rooms, kitchens, and bathrooms. Images are generated from COCO2017 [30] validation set annotations.
@@
Figure 7: This figure presents qualitative examples of closed-set configuration. Our model is capable of accurately rendering various classes of vehicles, such as cars, trains, motorcycles, and boats. Images are generated using grounding entities and image captions from the COCO2017 [30] validation set annotations.
@@
Figure 8: This figure presents qualitative examples of closed-set configuration. The portraits generated by ObjectDiffusion exhibit realistic facial expressions. Images are generated using grounding entities and image captions from the COCO2017 [30] validation set annotations.
@@
Figure 9: This figure represents qualitative examples of closed-set configuration. Our model demonstrates proficiency in generating high-resolution visuals of animals in their native habitat, achieving a very high AP precision score. Images are generated using grounding entities and image captions from COCO2017 [30] validation set annotations.
@@
Figure 10: This figure represents qualitative examples of closed-set configuration. It shows both static and dynamic scenes of different object categories. ObjectDiffusion can synthesize images of dishes as well as visuals of sports activities, such as surfing and skiing. Images are generated using grounding entities and image captions from COCO2017 [30] validation set annotations.
(a) Image Caption: “bright modern Scandinavian style house with a comfy couch, stylish coffee table, and sunlit windows homevibes relaxation”
Conditional Entities: a comfy couch, a stylish coffee table, a sunlit windows
(b) Image Caption: “a sleek, ultra-thin, high resolution bezel-less monitor, realistic, modern, detailed, vibrant colors, glossy finish, floating design, lowlight, art by peter mohrbacher and donato giancola, digital illustration, trending on Artstation, high-tech, smooth, minimalist workstation background, crisp reflection on screen, soft lighting”
Conditional Entities: a sleek, ultra-thin, high resolution bezel-less monitor, a keyboard, a mouse, a studying desk, a vase, a window
(c) Image Caption: “a waterfall and a modern high speed train running through the tunnel in a beautiful forest with fall foliage”
Conditional Entities: a waterfall, a modern high speed train
(d) Image Caption: “a beautiful digital painting depecting fluffy Pomeranian in a green forest, impressionism painting style, impressive, masterpeace”
Conditional Entities: a fluffy Pomeranian
Figure 11: This figure represents realistic and artistic qualitative examples of images generated in an open-set configuration with manual annotations. The images demonstrate the high quality and precision of ObjectDiffusion.